The formula R=-0.075t+3.85 can be used to predict the world record in the 1500 meter run, the years after 1930. Determine the inequality that identifies the years in which the world record will be less than 3.7 minutes. Solve for t>□ (Round to the nearest whole number.) The years after 1930, in which the record is less than 3.7 minutes, are described by the solution set {Years|Years>□}.
Answer by Theo(13342)

the equation is R = -.075 * T + 3.85

you want to know when R will be less than 3.7.

since R = -.075 * T + 3.85, then R < 3.7 when -.075 * T + 3.85 < 3.7

your inequality to solve is -.075 * T + 3.85 < 3.7

add .075 * T to both sides of this inequality and subtract 3.7 from both sides of this inequality to get:

3.85 - 3.7 < .075 * T

combine like terms to get:

.15 < .075 * T

divide both sides of this inequality by .075 to get:

.15 / .075 < T

solve for T to get T > .15 / .075 which results in T > 2

when T is greater than 2. the year is greater than 1930 + 2 = 1932.

the world record in the 1500 meter run will be less than 3.7 minutes after 1932.

R = -.0675 * T + 3.85 for T from 0 to 3 results in:

R = 0.0675 * 0 + 3.85 = 3.85 in 1930.
R = -.0675 * 1 + 3.85 = 3.7825 in 1931.
R = -.0675 * 2 + 3.85 = 1.7125 in 1932.
R = -.0675 * 3 + 3.85 = 3.6475 in 1933.